The pH of an aqueous solution of 0.296 M ammonium nitrate, NH NO3(aq), is : This solution is -
The options in the drop down menu read acidic/basic/neutral.
Answer #1

NH4NO3 dimociales completely in aqueous solution NHY NO3 Cagl | ) NH4 + (ap 1 + NO3- (aq) After disociation O 0.296M 01296M →5:56 X10-10 - 082 > x = 1.283 X10-5 M 0.296 Onu= [H3O+] = x = 1,283 X105M. bH = -luq [H30+] = 109 (1,283 X20-5) = I pH = 4.89